OpenStudy (anonymous):

At what point is the graph of y=1/2x^2 is the tangent line para llel to the line 2x-4y=3?

OpenStudy (raden):

first, m(slope) = y' = 1/2 * 2x = x see that the slope line 2x-4y=3 is m = -a/b (here a = 2, b = -4, and c = 3) so, m = -2/-4 = 1/2, back to first, get x = 1/2 then put x = 1/2 on the graph of y = 1/2 x^2, get y = 1/2 * (1/2)^2 = 1/8 therefore, the point is (x,y) = (1/2, 1/8)